OpenStudy (anonymous):

need help on 4x^3+9x^2+2x=0

OpenStudy (anonymous):

x(4x^2+9x+2) = 0 so either x = 0 or (4x^2+9x+2) = 0 4x^2+9x+2 = 0 => \[x = (-9\pm \sqrt{81-32})/8 => x = (-9 \pm 7)/8 => x = -1/4 or x = -2\]. So x = 0, or x = -1/4 or x = -2